1. The combustion of acetylene gas is represented by the equation: 2C2H2(g)+ 5 O2(g) 4CO2(g) + 2H2O (g) . How many grams of carbon dioxide (CO2) and grams of water(H2O)are produced when 66.0 g of C2H2 burns in oxygen? 2. If there are 35.o grams of C6H10 and 45.0 grams of O2, how many grams of the excess reagent will remain after the reaction ceases? 6 C6H10 +17 O2 12 CO2 + 10 H2O 3. For the balanced equation below, if the reaction of 77.0 grams of CaCN2 produces 27.1 g of NH3, what is the percent yield? CaCN2 + 3 H2O CaCO3 + 2 NH3. I need an explanation to your answers everyone. Thank you so much.


1
Expert's answer
2021-01-13T13:55:24-0500

1. M(C2H2) = 26 g/mol

m(C2H2) "= \frac{66.0}{26}=2.538 \;mol"

n(CO2) = 2n(C2H2) "= 2 \times 2.538 = 5.076 \;mol"

M(CO2) = 44 g/mol

m(CO2) "= 44 \times 5.076 = 223.34 \;g"

n(H2O) = n(C2H2) = 2.538 mol

M(H2O) = 18 g/mol

m(H2O) "= 18 \times 2.538 = 45.68 \;g"

2. M(C6H10) = 82 g/mol

n(C6H10) "= \frac{35.0}{82}=0.426 \;mol"

M(O2) = 32 g/mol

n(O2) "= \frac{45.0}{32}=1.406 \;mol"

Theoretical proportion:

C6H10 : O2 = 6 : 17 = 1 : 2.83

Real proportion:

C6H10 : O2 = 0.426 : 1.406 = 1 : 3.3

So, C6H10 is the limiting reactant.

According to the reaction:

n(O2) "= \frac{17}{6}n(C_6H_{10}) = \frac{17}{6} \times 0.426 = 1.207 \;mol"

Used mass of O2:

m’(O2) "= 1.207 \times 32 = 38.624 \;g"

Δm(O2) = 45.0 – 38.624 = 6.376 g

So, 6.376 g of O2 (excess reagent) will remain after the reaction ceases.

3. M(CaCN2) = 80 g/mol

n(CaCN2) "= \frac{77.0}{80}= 0.9625 \;mol"

According to the reaction:

n(NH3) = 2n(CaCN2) "= 2 \times 0.9625 = 1.925 \;mol"

M(NH3) = 17 g/mol

m’(NH3) = "1.925 \times 17" = 32.725 g (theoretical mass)

Proportion:

32.725 – 100 %

27.1 – x

"x =\frac{27.1 \times 100}{32.725} = 82.81\; \%"

So, he percent yield is 82.81 %.